# Enhancing System-Level Safety in Mixed-Autonomy Platoon via Safe Reinforcement Learning

### 20 Jan 2024 v1

@@ IV-C1 CBF-QP-Based Approach to Incorporate Safety Guarantees
where $\tau$ is the minimum allowed time headway. CAV is indexed by $i$ and the following HDVs are indexed by $i+1,\cdots,n$. Using the time headway as the CBF candidate, we have:
@@
However, for the CBF candidate $h_{\text{th},j}(x)$, the relative degree [^2] The relative degree represents the number of times we need to differentiate the CBF candidate along the dynamics of Eq. (9) until the control action $u$ explicitly shows. is $j-i+1\geq 2$ for the $j$-th following vehicle, for which the original CBF can not handle. Although such issues can be handled by high-order CBFs [4], obtaining the high-order derivatives of states in real-world systems is typically computationally expensive and imprecise. Therefore, we introduce an alternative approach to reduce the relative degrees. Specifically, instead of directly using time headway as the CBF candidate, we utilize the revised safety conditions as:
$h_{i}(x)$ $=h_{\text{th},i}(x)=s_{i}-\tau v_{i}$ (20)
@@
$\quad\quad-\tau\hat{F}_{j}\left(x_{j},v_{j-1}\right),j\in\left\{i+1,\cdots,n\right\}$ (21b)
$L_{g}h_{i}=\nabla h_{i}(x)^{T}g(x)=-\tau$ (21c)
@@
$L_{f}h_{i}+L_{g}h_{i}u+\alpha_{\text{CBF},i}(h_{i})\geq 0$ (22)
$L_{f}h_{j}+L_{g}h_{j}u+\alpha_{\text{CBF},j}(h_{j})+\sigma_{j}\geq 0,j\in\left\{i+1,\cdots,n\right\}$
where the class $\mathcal{K}$ function $\alpha_{\text{CBF},i}(\cdot),\cdots,\alpha_{\text{CBF},n}(\cdot)$ are set to linear functions, for simplicity, with positive coefficients $k_{i},\cdots,k_{n}$ serving as the parameters to be trained. $\sigma_{i+1}\cdots\sigma_{n}$ are the slack variables for HDV safety constraints to avoid conflict with CAV safety constraints. To better tailor the derived CBF constraint to the DRL settings, we rewrite (22) as follows:
$L_{f}h_{i}+L_{g}h_{i}(u_{\text{safe}}+u_{\text{RL}})+k_{i}h_{i}\geq 0$ (23a)
@@
where $u=u_{\text{RL}}+u_{\text{safe}}$ with $u_{\text{RL}}$ representing the action provided by the DRL algorithm. We divide the control input into $u_{\text{RL}}$ and $u_{\text{safe}}$ to highlight the compensation effect of the CBF-QP.
@@
$a_{\min}\leq u_{\text{RL}}+u_{\text{safe}}\leq a_{\max}.$ (24)
where $a_{\max}$ and $a_{\min}$ denote the maximum and minimum acceleration rates for CAVs, respectively. However, the acceleration constraint may conflict with the CAV safety constraint 23a and make the QP problem infeasible, which is because the CAV safety constraint is a hard constraint without a relaxation term. To guarantee the feasibility of the QP problem, we integrate a feasibility constraint as in [33, 34]. By utilizing this approach, we enforce a sufficient condition for the feasibility via another CBF in the mixed-autonomy platoon. The feasibility constraint is formulated as follows:
$u_{\text{safe}}+u_{\text{RL}}\leq\hat{F}_{i-1}+k_{\text{f}}(v_{i-1}-v_{i}-\tau a_{\min})$ (25)
with a positive trainable parameter $k_{\text{f}}$ and estimated acceleration of the preceding vehicle $\hat{F}_{i-1}$. The feasibility constraint provides a conflict-free guarantee of the QP problem corresponding to the next time interval, the details of which can be found in [34].
@@
s.t. $L_{\hat{f}}h_{i}+L_{g}h_{i}(u_{\text{safe}}+u_{\text{RL}})+k_{i}h_{i}\geq 0,$
$L_{f}h_{j}+L_{g}h_{j}(u_{\text{safe}}+u_{\text{RL}})+k_{j}h_{j}+\sigma_{j}\geq 0,$
@@
$u_{\text{safe}}+u_{\text{RL}}\leq\hat{F}_{i-1}+k_{\text{f}}(v_{i-1}-v_{i}-\tau a_{\min}),$
@@
where $b_{j}>0,~{}j=i+1,\cdots,n$ is the penalty coefficient for the slack variable $\sigma_{j}$. Here, $b_{j}>0,~{}j=i+1,\cdots,n$ are all set to $1$ for simplicity. Parameters $\{k_{j}\}_{j=i}^{n}$ and $k_{\text{f}}$ can be trained simultaneously with the RL policy network via a differentiable safety layer, which allows the CBF constraints to better adapt to the specific environment, as presented next.
@@ IV-C2 Differentiable QP for Neural Networks
We assume that the loss function output is $\ell$, and the optimal solution of the CBF-QP is $w^{\star}=\left(u_{\text{safe}},\sigma_{1},\cdots,\sigma_{n}\right)$, so the derivative of loss from the previous backward pass can be written as $\frac{\partial\ell}{\partial w^{\star}}$. To train the RL module and the safety layer, we are interested in calculating the following partial derivative:
$\frac{\partial\ell}{\partial\theta}=\frac{\partial\ell}{\partial w^{\star}}\frac{\partial w^{\star}}{\partial u_{\text{RL}}}\frac{\partial u_{\text{RL}}}{\partial\theta}$ (27)
Since $\frac{\partial\ell}{\partial w^{\star}}$ can be easily obtained from the form of the loss function, and $\frac{\partial u_{\text{RL}}}{\partial\theta}$ is defined by the architecture of the RL embedded neural networks and the trainable parameters of the safety layer, which are parameterized by $\theta=\left\{\theta_{\text{RL}},\{k_{i}\}_{i=0}^{n},k_{f}\right\}$, we only need to calculate $\frac{\partial w^{\star}}{\partial u_{\text{RL}}}$, as presented below.
@@
subject to $Gw\leq q(u_{\text{RL}})$
where $Q,G,q$ are the corresponding parameters of the QP problem. Note that only $q$ is a function of $u_{\text{RL}}$. As in [23], the calculation of $\frac{\partial w^{\star}}{\partial u_{\text{RL}}}$ can be performed by differentiating the KKT conditions with equality, i.e.,
$Qw^{\star}+G^{T}\lambda^{\star}$ $=0$ (29)
$D\left(\lambda^{\star}\right)\left(Gw^{\star}-q(u_{\text{RL}})\right)$ $=0,$
where $w^{\star}$ and $\lambda^{\star}$ represent the optimal primal and dual variables, and $D(\lambda^{\star})$ is a diagonal matrix constructed from vector $\lambda^{\star}$.
Then we take the derivative of the RL action $u_{\text{RL}}$ and write in the matrix form as:
${\left[\begin{array}[]{l}\displaystyle\frac{\partial w^{\star}}{\partial u_{\text{RL}}}\\ \displaystyle\frac{\partial\lambda^{\star}}{\partial u_{\text{RL}}}\end{array}\right]=}K^{-1}\left[\begin{array}[]{c}\mathbb{O}\\ D\left(\lambda^{\star}\right)\displaystyle\frac{\partial q(u_{\text{RL}})}{\partial u_{\text{RL}}}\end{array}\right].$ (30)
with $K=\left[\begin{array}[]{cc}Q&G^{T}\\ D\left(\lambda^{\star}\right)G&D\left(Gw^{\star}-q(u_{\text{RL}})\right)\end{array}\right]$.
@@ V Simulation Results
In this section, we evaluate the proposed method in simulations of a typical mixed-autonomy platoon. We consider a typical platooning scenario as depicted in Fig. 1, which features a platoon of five vehicles traveling on a single lane, with the third vehicle being a CAV and the others HDVs.
@@ V-A Training Settings and Results
In the training scenario, we utilize a random velocity disturbance setup, wherein the leading vehicle’s velocity disturbance is sampled from a Gaussian distribution with zero mean and a standard deviation of $2$ m/s independently at each time step within an episode. This chosen velocity disturbance is then multiplied by the time step ($0.1$s) and added to the head vehicle’s original velocity. For the parameters for the HDV car-following model (e.g., OVM [28]), we consider $\alpha=0.6,\beta=0.9,s_{\text{st}}=5,s_{\text{go}}=35$. The equilibrium spacing and velocity are $20$ m and $15$ m/s.
The training hyperparameters are as follows. The learning rates for the actor and critic are both set at 0.0003 and we employ a linear decay schedule for the learning rate to facilitate training. As such, the performance of the algorithms is not sensitive to the initial learning rate and we choose the ones that can make the training process converge rapidly. The batch size is $2048$, the decay factor of the advantage function is $0.95$, the PPO clip parameter is set to $0.2$, and the number of experience utilizing is set to $10$. The training episodes are $500$. For the CBF parameters, time headway $\tau$ is set to $0.3$s and the initial values for the trainable parameters $\left\{k_{0},k_{1},k_{2},k_{f}\right\}$ are $\left\{1,1,1,10\right\}$.
Under the specified training scenarios and parameter settings, the training rewards per episode for PPO and PPO with the safety layer are depicted in Fig. 3. Both algorithms converge after a sufficient number of training episodes, indicating that the designed reward function is suitable for mixed-autonomy platoon environments. Notably, the algorithms with a safety layer converge faster. This is expected as the safety layer reduces the range of feasible parameters and thus can accelerate training, especially in scenarios with large parameter space.
For the setting of the learning-based driver behavior identification as in Eq. (16), the online training for the two networks $\eta_{\xi,i}$ and $\zeta_{\psi,i}$ are separated. $\eta_{\xi,i}$ is training to estimate the linearized car-following model output, then $\zeta_{\psi,i}$ is training to estimate the bias of the true car-following model and linearized car-following model. The learning rates for both networks are 0.0001. To illustrate the benefits of our proposed learning-based driver behavior identification method, we conduct a comparison with Recursive Least Square (RLS) in random velocity disturbance setup for the estimation accuracy of vehicle $4$. The mean square error is 0.00216 for the proposed method and 0.00251 for RLS. The comparison results are given in Fig. 4, which shows that our proposed method yields better accuracy for acceleration rate prediction compared with RLS.

### 01 Mar 2024 v2

@@ IV-C1 CBF-QP-Based Approach to Incorporate Safety Guarantees
where $\tau$ is the minimum allowed time headway. CAV is indexed by $i$ and the following HDVs are indexed by $i+1,\cdots,n$. With Eq. (15), we have the following CBF candidate:
@@
We make the following remark for the CBF candidate Eq. (16), which can also be seen as a state constraint.
Remark 1
Although alternative methods exist to handle Eq. (16), it is challenging for these methods to provide a formal safety guarantee in a computationally efficient manner. One common method is to integrate Eq. (16) into the DRL reward function, which, however, is only a soft constraint without a formal guarantee. Alternatively, Eq. (16) can be explicitly integrated into an optimization problem to serve as a state constraint, which can be inefficient to account for future states due to the myopic nature of the state constraint (i.e., it only involves a particular time step). In contrast, CBF as introduced in Def. 1 can provide a theoretical safety guarantee in a computationally efficient manner via the property of forward invariance, which ensures that if the current state is safe, there always exist control actions such that future states remain safe.
However, for the CBF candidate $h_{\text{th},j}(x)$, the relative degree [^2] The relative degree represents the number of times we need to differentiate the CBF candidate along the dynamics of Eq. (7) until the control action $u$ explicitly shows. is $j-i+1\geq 2$ for the $j$-th following vehicle, which cannot be handled by the original CBF. Although such an issue can be handled by high-order CBFs [4], obtaining the high-order derivatives of states in real-world systems is typically computationally expensive and imprecise. Therefore, we introduce an alternative approach to reduce the relative degrees. Specifically, instead of directly using Eq. (17) as the CBF candidate, we utilize revised safety conditions:
$h_{i}(x)$ $=h_{\text{th},i}(x)=s_{i}-\tau v_{i},$ (17)
@@
$\quad\quad-\tau\hat{\mathbb{F}}_{j}\left(x_{j},v_{j-1}\right),j\in\left\{i+1,\cdots,n\right\},$ (18b)
$L_{g}h_{i}=\nabla h_{i}(x)^{T}g(x)=-\tau,$ (18c)
@@
$L_{f}h_{i}+L_{g}h_{i}u+\alpha_{\text{CBF},i}(h_{i})\geq 0,$ (19a)
$L_{f}h_{j}+L_{g}h_{j}u+\alpha_{\text{CBF},j}(h_{j})+\sigma_{j}\geq 0,j\in\left\{i+1,\cdots,n\right\}$ (19b)
where the class $\mathcal{K}_{\infty}$ function $\alpha_{\text{CBF},i}(\cdot),\cdots,\alpha_{\text{CBF},n}(\cdot)$ are set to linear functions with positive coefficients $k_{i},\cdots,k_{n}$ serving as the parameters to be trained. The choice of a linear class $\mathcal{K}_{\infty}$ function follows the insights of [32], which showed that using a linear class $\mathcal{K}_{\infty}$ function for CBFs, compared to a quadratic one, results in smaller, more gradual control inputs that can respect CAV actuation limits and enhance passenger comfort. $\sigma_{i+1}\cdots\sigma_{n}$ are the slack variables for HDV safety constraints to avoid conflict with CAV safety constraints. To better tailor the derived CBF constraint to the DRL settings, we rewrite Eq. (IV-C1) as follows:
$L_{f}h_{i}+L_{g}h_{i}(u_{\text{safe}}+u_{\text{RL}})+k_{i}h_{i}\geq 0,$ (20a)
@@
where $u=u_{\text{RL}}+u_{\text{safe}}$ with $u_{\text{RL}}$ representing the action provided by the DRL algorithm. We divide the control input into $u_{\text{RL}}$ and $u_{\text{safe}}$ to highlight the compensation effect of the CBF-QP. Notice that directly optimizing $u$ is equivalent to optimizing $u_{\text{safe}}$, as $u_{\text{RL}}$ is treated as given in the QP problem.
@@
$a_{\min}\leq u_{\text{RL}}+u_{\text{safe}}\leq a_{\max},$ (21)
where $a_{\max}$ and $a_{\min}$ denote the maximum and minimum acceleration rates for CAVs, respectively. However, the acceleration constraint may conflict with the CAV safety constraint Eq. (20a) and make the QP problem infeasible, which is because the CAV safety constraint is a hard constraint without a relaxation term. To guarantee the feasibility of the QP problem, we integrate a feasibility constraint as in [40, 41]. By utilizing this approach, we enforce a sufficient condition for the feasibility via another CBF in the mixed-autonomy platoon. The feasibility constraint is formulated as follows:
$u_{\text{safe}}+u_{\text{RL}}\leq\hat{\mathbb{F}}_{i-1}+k_{\text{f}}(v_{i-1}-v_{i}-\tau a_{\min}),$ (22)
with a positive trainable parameter $k_{\text{f}}$ and estimated acceleration of the preceding vehicle $\hat{\mathbb{F}}_{i-1}$. The feasibility constraint provides a conflict-free guarantee of the QP problem corresponding to the next time interval, the details of which can be found in [41].
@@
s.t. $L_{f}h_{i}+L_{g}h_{i}(u_{\text{safe}}+u_{\text{RL}})+k_{i}h_{i}\geq 0,$
$L_{\hat{\mathbb{F}}_{j}}h_{j}+L_{g}h_{j}(u_{\text{safe}}+u_{\text{RL}})+k_{j}h_{j}+\sigma_{j}\geq 0,$
@@
$u_{\text{safe}}+u_{\text{RL}}\leq\hat{\mathbb{F}}_{i-1}+k_{\text{f}}(v_{i-1}-v_{i}-\tau a_{\min}),$
@@
where $b_{j}>0,~{}j=i+1,\cdots,n$ is the penalty coefficient for the slack variable $\sigma_{j}$. Here, $b_{j}>0,~{}j=i+1,\cdots,n$ are all set to $1$ for simplicity. Parameters $\theta_{\text{CBF}}=\left\{\{k_{j}\}_{j=i}^{n},k_{\text{f}}\right\}$ can be trained simultaneously with the RL policy network via a differentiable QP layer, which allows the CBF constraints to better adapt to the specific environment, as presented next.
@@ IV-C2 Differentiable QP for Neural Networks
Let us denote the loss function by $\ell$, and the optimal solution of the CBF-QP is $w^{\star}=\left(u_{\text{safe}},\sigma_{1},\cdots,\sigma_{n}\right)$. Let $\theta=\left\{\theta_{\text{RL}},\theta_{\text{CBF}}\right\}$ represent the trainable parameters in the DRL actor network and the CBF-QP controller. To train these parameters, we are interested in calculating the following partial derivatives, which are obtained based on the chain rule:
$\frac{\partial\ell}{\partial\theta_{\text{RL}}}$ $=\frac{\partial\ell}{\partial u}\frac{\partial u}{\partial\theta_{\text{RL}}}=\frac{\partial\ell}{\partial u}\Big{(}\frac{\partial u_{\text{RL}}}{\partial\theta_{\text{RL}}}+\frac{\partial u_{\text{safe}}}{\partial\theta_{\text{RL}}}\Big{)}$
$=\frac{\partial\ell}{\partial u}\Big{(}\frac{\partial u_{\text{RL}}}{\partial\theta_{\text{RL}}}+\frac{\partial u_{\text{safe}}}{\partial u_{\text{RL}}}\frac{\partial u_{\text{RL}}}{\partial\theta_{\text{RL}}}+\frac{\partial u_{\text{safe}}}{\partial\theta_{\text{CBF}}}\frac{\partial\theta_{\text{CBF}}}{\partial\theta_{\text{RL}}}\Big{)}$
$=\frac{\partial\ell}{\partial u}\Big{(}\frac{\partial u_{\text{RL}}}{\partial\theta_{\text{RL}}}+\frac{\partial u_{\text{safe}}}{\partial u_{\text{RL}}}\frac{\partial u_{\text{RL}}}{\partial\theta_{\text{RL}}}\Big{)},$ (24)
$\frac{\partial\ell}{\partial\theta_{\text{CBF}}}$ $=\frac{\partial\ell}{\partial u}\frac{\partial u}{\partial\theta_{\text{CBF}}}=\frac{\partial\ell}{\partial u}\left(\frac{\partial u_{\text{RL}}}{\partial\theta_{\text{CBF}}}+\frac{\partial u_{\text{safe}}}{\partial\theta_{\text{CBF}}}\right)=\frac{\partial\ell}{\partial u}\frac{\partial u_{\text{safe}}}{\partial\theta_{\text{CBF}}}$ (25)
where the first equality condition in both Eq. (24) and Eq. (25) results from the chain rule, and the second equality condition in both equations results from the relation $u=u_{\text{RL}}+u_{\text{safe}}$. The third equality condition in Eq. (24) is because by Eq. (23), $u_{\text{safe}}$ can be seen as a function with respect to $u_{\text{RL}}$ and $\theta_{\text{CBF}}$. The last equality condition of Eq. (24) is due to the relation $\frac{\partial\theta_{\text{CBF}}}{\partial\theta_{\text{RL}}}=0$ because the CBF-QP parameters $\theta_{\text{CBF}}$ do not depend on the actor network. The last equality condition of Eq. (25) is due to the relation $\frac{\partial u_{\text{RL}}}{\partial\theta_{\text{CBF}}}=0$ because the actor network does not depend on the parameters $\theta_{\text{CBF}}$. Note that $\frac{\partial\ell}{\partial u}$ can be easily obtained from the loss function, and $\frac{\partial u_{\text{RL}}}{\partial\theta}$ is defined by the architecture of the actor network. Therefore, we only need to calculate $\frac{\partial u_{\text{safe}}}{\partial u_{\text{RL}}}$ and $\frac{\partial u_{\text{safe}}}{\partial\theta_{\text{CBF}}}$, which are part of $\frac{\partial w^{\star}}{\partial u_{\text{RL}}}$ and $\frac{\partial w^{\star}}{\partial\theta_{\text{CBF}}}$ as presented next.
@@
subject to $Gw\leq q(u_{\text{RL}},\theta_{\text{CBF}})$
where $Q,G,q$ are the corresponding parameters of the QP problem. Note that only $q$ is a function of $u_{\text{RL}}$ and $\theta_{\text{CBF}}$. As in [30], the calculation of $\frac{\partial w^{\star}}{\partial u_{\text{RL}}}$ and $\frac{\partial w^{\star}}{\partial\theta_{\text{CBF}}}$ can be performed by differentiating the KKT conditions with equality, i.e.,
$Qw^{\star}+G^{T}\lambda^{\star}$ $=0,$ (27)
$\mathbb{D}\left(\lambda^{\star}\right)\left(Gw^{\star}-q(u_{\text{RL}},\theta_{\text{CBF}})\right)$ $=0,$
where $w^{\star}$ and $\lambda^{\star}$ represent the optimal primal and dual variables, and $\mathbb{D}(\lambda^{\star})$ is a diagonal matrix constructed from vector $\lambda^{\star}$.
Since the calculation of $\frac{\partial u_{\text{safe}}}{\partial u_{\text{RL}}}$ and $\frac{\partial u_{\text{safe}}}{\partial\theta_{\text{CBF}}}$ is similar, we next only illustrate the calculation of $\frac{\partial u_{\text{safe}}}{\partial u_{\text{RL}}}$. For the calcluation of $\frac{\partial w^{\star}}{\partial\theta_{\text{CBF}}}$, we only need to replace $u_{\text{RL}}$ with $\theta_{\text{CBF}}$ in Eq.(28). Then we take the derivative of Eq.(27) with respect to the RL action $u_{\text{RL}}$ and write in the matrix form as:
${\left[\begin{array}[]{l}\displaystyle\frac{\partial w^{\star}}{\partial u_{\text{RL}}}\\ \displaystyle\frac{\partial\lambda^{\star}}{\partial u_{\text{RL}}}\end{array}\right]=}K^{-1}\left[\begin{array}[]{c}\mathbb{O}\\ \mathbb{D}\left(\lambda^{\star}\right)\displaystyle\frac{\partial q(u_{\text{RL}},\theta_{\text{CBF}})}{\partial u_{\text{RL}}}\end{array}\right],$ (28)
with $K=\left[\begin{array}[]{cc}Q&G^{T}\\ \mathbb{D}\left(\lambda^{\star}\right)G&\mathbb{D}\left(Gw^{\star}-q(u_{\text{RL}},\theta_{\text{CBF}})\right)\end{array}\right]$.
@@ V Simulation Results
In this section, we evaluate the proposed method in simulations of a typical mixed-autonomy platoon. We consider a typical platooning scenario as depicted in Fig. 1, which features a platoon of five vehicles traveling on a single lane, with the third vehicle being a CAV and the others HDVs. For benchmarks, we compare four types of models: pure car following, PPO without the safety layer (PPO without safety guarantee), PPO with the safety layer but without learning-based human driver behavior identification (i.e., safe-RL without SI), and PPO with the safety layer and learning-based human driver behavior identification (i.e., safe-RL with SI).
@@ V-A Training Settings and Results
In the training scenario, we utilize a random velocity disturbance setup, wherein the head vehicle’s velocity disturbance is sampled from a Gaussian distribution with zero mean and a standard deviation of $2$ m/s independently at each time step within an episode. This chosen velocity disturbance is then multiplied by the time step ($0.1$ s) and added to the head vehicle’s original velocity.
Without loss of generality, we choose the optimal velocity model (OVM) [42] as the car-following model $\mathbb{F}$, following the setting of LCC [7]. Note that the OVM can be replaced by any car-following model without influencing the applicability of the proposed method. The OVM reads as follows:
The parameters of OVM is set as $\alpha=0.6,\beta=0.9,s_{\text{st}}=5,s_{\text{go}}=35$. The equilibrium spacing and velocity are $20$ m and $15$ m/s.
The training hyperparameters are as follows. The learning rates for the actor and critic networks are both set at 0.0003 and we employ a linear decay schedule for the learning rate to facilitate training. As such, the performance of the algorithms is not sensitive to the initial learning rate and we choose the ones that can make the training process converge rapidly. The batch size is $2048$, the decay factor of the advantage function is $0.95$, the PPO clip parameter is set to $0.2$, and the number of experience utilizing is set to $10$. The training episodes are $500$. For the CBF parameters, the minimum time headway $\tau$ is set to $0.3$s and the initial values for the trainable parameters $\left\{k_{0},k_{1},k_{2},k_{f}\right\}$ are $\left\{1,1,1,10\right\}$.
Under the specified training scenarios and parameter settings, the training rewards per episode for PPO without safety guarantee (i.e., without safety layer) and safe-RL with SI (i.e., with safety layer) are depicted in Fig. 3. Both algorithms converge after a sufficient number of training episodes, indicating that the designed reward function is suitable for mixed-autonomy platoon environments. Notably, the algorithms with a safety layer converge faster. This is expected as the safety layer reduces the range of feasible parameters and thus can accelerate training, especially in scenarios with large parameter space.
For the learning-based driver behavior identification module as in Eq. (14), we train the parameters $\eta_{\xi,i}$ for the computation graph and $\zeta_{\psi,i}$ for the neural network-based bias estimator. The learning rate is set to 0.0001. To illustrate the benefits of our proposed learning-based driver behavior identification method, we conduct a comparison with recursive least squares (RLS) in terms of the estimation accuracy of vehicle $4$. The mean square error is 0.0019 for the proposed method and 0.0024 for RLS. The comparison results are given in Fig. 4, which shows that our proposed method yields better accuracy for acceleration rate prediction compared with RLS.